Next number in sequence for the following:
7 16 8 27 9
3 -6 12 4 20
2 7 26 101 400
7 21 8 72 9

Answer by Edwin McCravy(20077)

7 16 8 27 9
The odd numbered terms go 7, 8, 9, 11
The even numbered terms 8×2, 9×3.  Therefore
The next term is 10×4 or 40 <-- the next number.

3 -6 12 4 20
Start with 3, 
subtract 9, get -6
add twice 9, which is 18, get 12
subtract 8, get 4
add twice 8, which is 16, get 20.  Therefore,
subtract 7, get 13  <-- the next number.

2 7 26 101 400
Start with 2,
2×4=8, 8-1=7
7×4=28, 28-2=26
26×4=104, 104-3 = 101
101×4=404, 404-4 = 400.  Therefore,
400×4=1600, 1600-5 = 1595 <-- the next number.

7 21 8 72 9
The odd numbered terms go 7, 8, 9, 11, like the first sequence above
The even numbered terms go 7×3=21, 8×3×3=72, Therefore
the next term is 9×3×3×3 = 243 <-- the next number.

Edwin